# Understanding Addition In Transformers

### 29 Feb 2024 v6

@@ 8 Conclusions and Future Work
Our theoretical framework of necessary addition subtasks provides a foundation for the model’s behavior. The digit-wise training loss curves reveal independent refinement consistent with separate digit-specific circuits. Attention patterns illustrate staging and time-ordering of operations. Controlled ablation experiments validate ouur hypothesis about algorithmic elements’ roles. Together these methods enable a detailed accounting of the model’s addition procedure.
@@
Our study paves the way for numerous potential research avenues. Recognizing the challenges in rare cases can inspire methods to enhance the robustness of addition models. The established framework might be adapted to elucidate models for integer subtraction or multiplication. By integrating proven and effective addition modules into a larger, untrained network geared towards multiplication, the training process could be expedited. Further, decoding the multiplication algorithm becomes more straightforward when the addition-related tasks are already recognized and deemed reliable. Utilizing this modular approach can simplify the understanding of intricate algorithms, propelling advancements in the field of mechanistic interpretability. In summary, this research underscores that diving deep into the workings of contemporary machine learning can highlight valuable strengths, pinpoint areas for improvement, and present avenues for accelerated progress.
@@ 9 Reproducibility Statement
To ensure our work is reproducible, we provide the full source code in the supplementary materials, as well as all necessary data and parameters and instructions to reproduce our experimental results.
@@ Appendix A Appendix - Number of Attention Heads
The model can be successfully trained with 2, 3 or 4 attention heads. However as described in (Fig. 10), 3 attention heads is more useful for human understanding than 2 or 4 heads.
@@
Figure 11: For 5 digit addition, using 2 attention heads works, but the model attends to multiple token pairs in a single head, suggesting that multiple tasks are being packed into a single head, which makes it harder to understand.
@@ Appendix B Appendix - Training Loss by Number of Digits
The model can be successfully trained with 2, 5, 10, 15, etc digits. As expected, for a given loss threshhold, 15 digit addition takes longer to train than 10 digit addition, which takes longer to train than 5 digit addition.
Figure 12: The per-digit training loss curves for 5 digit addition
Figure 13: The per-digit training loss curves for 10 digit addition.
Figure 14: The per-digit training loss curves for 15 digit addition.
@@ Appendix C Appendix - Attention Patterns by Number of Digits
The model can be successfully trained with 2, 5, 10, 15, etc digits. The attention patterns for theses cases show similarities that aid in human understanding.
Figure 15: For 5 digit addition, and 3 attention heads, the attention pattern has a strong double-staircase shape. Each step of the staircase is 3 blocks wide, showing the heads are attending to different tokens in each row.
Figure 16: For 10 digit addition, and 3 attention heads, the attention pattern has a strong double-staircase shape. Each step of the staircase is 3 blocks wide, showing the heads are attending to different tokens in each row.
Figure 17: For 15 digit addition, and 3 attention heads, the attention pattern has a strong double-staircase shape. Each step of the staircase is 3 blocks wide, showing the heads are attending to different tokens in each row.
@@ Appendix D Appendix - Model Configuration
Experimentation was done in a CoLab notebook:
  • It will be loaded to github post publication.
@@
  • The key parameters (which can all be experimented with) are:
@@
  • During a training run the model processes about 1.5 million training datums. For the 5 digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So the training data is much less than 1% of the possible problems.
  • Because US9 cascades (e.g. 44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare, the data generator was enhanced to increase the likelihood of these cases turning up in the training data.
The CoLab notebook can be downloaded from https://github.com/apartresearch/conceptual-interp
@@ Appendix E Appendix - Solving a jigsaw using many people
Here is useful analogy for changing how we approach a problem to more closely resemble how a Transformer would; a jigsaw puzzle. A single person could solve it using a combination of meta knowledge of the problem (placing edge pieces first), categorisation of resources (putting like-coloured pieces into piles), and an understanding of the expected outcome (looking at the picture on the box).
@@ Appendix F Appendix - Use Sum 9 Training Loss Graph
Fig. 18 shows a training loss curve for just BA and UC1 . The model has high loss on Use Sum 9 cascades shown as high variability in the loss graph (See Fig. 19).
Figure 18: Training loss for digit 3 showing that, in Phase 2, the refining of Use Carry 1 lags behind Base Add . This supports the claim that Base Add and Use Carry 1 are refined separately and have separate calculation algorithms.
Figure 19: Training loss for digit 3 for the Base Add , Use Carry 1 and Use Sum 9 tasks, showing the model struggles to reduce loss on Use Sum 9 compared to Base Add and Use Carry 1 .
@@ Appendix G Appendix - Model Algorithm as Pseudocode
This section reproduces the algorithm summarised in Fig. 8 as pseudocode. This code calculates the largest digit first, handling simple (but not cascading) UseSum9 cases, and returning the answer.
@@
3:   for all $i=0,\dots,n\ digits-1$ do $\triangleright$ Loop over each digit
4:      $\text{pos}\leftarrow n\ digits-i-1$triangleright Current position from the right
@@
9:         if $q1_{\text{prev prev pos}}+q2_{\text{prev prev pos}}\geq 10$ then
@@
15:         if $q1_{\text{prev pos}}+q2_{\text{prev pos}}\geq 10$ then
@@
21:         if $q1_{\text{prev pos}}+q2_{\text{prev pos}}==9$ then
@@
34:      $\text{digitanswer}\leftarrow\textsc{MODULUS}(\text{digitanswer},10)$triangleright Correct the current digit if it’s $\geq 10$
35:      $\text{answer}\leftarrow\text{digitanswer}+\text{answer}\times 10$triangleright Concatenate the current digit to the answer
36:   end forreturn answer $\triangleright$ Return the final calculated answer
37:end function
@@ Appendix H Appendix - The Loss function and loss measures
The final training loss varies with the number of digits in the question as shown in Tab. 1.
@@
The final training loss for each digit in the question varies as shown in Tab. 2.
@@
Table 2: For 5-digit addition, final training loss per digit.
Using our mathematics framework, we categorize each training question as one of following:
  • BA : BaseAdd questions only require “base add” calculations to get the correct answer.
  • MC1 : MakeCarry1 questions require “use carry 1” and “base add” calculations to get the correct answer.
  • US9 : UseSum9 questions require “use sum 9”, “use carry 1” and “base add” calculations to get the correct answer.
This allows us to measure the loss per category as shown in Tab. 3. The frequency of these question types differ significantly ( BA = 61%, MC1 =33%, US9 =6% ) so the final training loss values, at this level, are not very informative.
@@
After training, we used the model to give answers to questions. To understand whether the calculations at token n are important, we look at the impact on loss of ablating all attention heads at that token (using the TransformerLens framework, zero ablation of the blocks.0.hook_resid_post data set, the above loss function, and a “cut off” loss threshold of 0.08). Tab. 4 shows sample results.
@@
Table 4: For 5 digit addition, for the 18 tokens / calculation rows, how abalting each token/row impacts the calculation loss.
This shows that all important calculations are completed while the model processes just 6 of the 18 tokens.
For deeper investigation, we created 1̃00 hand-curated test questions. They cover all question types (BA, MC1 and MS9) and all the answer digits (A5 .. A0) that these question types can occur in. These test cases were not used in model training. Example questions include:
@@
 11 ’Nyyyyy’: 47, ’NyNyyy’: 5, ’yNyyyy’: 7, ’yyNyyy’: 7, ’yyyNyy’: 1, …
12 ’yNyyyy’: 97, ’yNNyyy’: 7, ’NNyyyy’: 3, ’Nyyyyy’: 2, ’yyyNyy’: 1, …
13 ’yyNyyy’: 85, ’NyNyyy’: 3, ’yNNyyy’: 2, ’Nyyyyy’: 3, ’yNyyyy’: 5, …
14 ’yyyNyy’: 72, ’yyNNyy’: 4, ’NyyNyy’: 3, ’yNyNyy’: 3, ’Nyyyyy’: 2, …
15 ’yyyyNy’: 74, ’yyNyNy’: 3, ’Nyyyyy’: 2, ’NyyyNy’: 3, ’yNyyNy’: 3, …
16 ’yyyyyN’: 82, ’yyNyyN’: 2, ’NyyyyN’: 4, ’yNyyyN’: 4, ’yyNyyy’: 9, …
Table 5: For 5 digit addition, how ablating all the attention heads for a row impacts answer digit correctness. An ‘N’ means that answer digit was incorrect in the predicted answer. In each step, the failure count for the top grouping is > 6 times the next most common grouping - evidence that each step calculates one answer digit. There is a very steady progression through the rows with each successive row focusing on the next answer digit - this progression reflects numerically the “double staircase” attention pattern.

### 18 Mar 2024 v8

@@ 9 Conclusions
Our theoretical framework of necessary addition subtasks provides a foundation for the model’s behavior. The digit-wise training loss curves reveal independent refinement consistent with separate digit-specific circuits. Attention patterns illustrate staging and time-ordering of operations. Controlled ablation experiments validate our hypothesis about algorithmic elements’ roles. Together these methods enable a detailed accounting of the model’s addition procedure.
@@
10 Limitations and Future Work
One concrete limitation of our current model is its difficulty handling the rare case of adding two 9-digit numbers that sum to 1 billion (e.g. 999,999,999 + 1). Further investigation is needed to understand why this specific edge case poses a challenge and to develop strategies to improve the model’s performance, such as targeted data augmentation or architectural modifications.
Regarding the MLP component, a detailed ablation study could elucidate its precise role and contributions. Systematically removing or retraining this component while monitoring performance changes could shed light on whether it is essential for the overall algorithm or plays a more auxiliary part. For future work, a natural next step is to apply our framework to reverse-engineer integer subtraction models. Subtraction shares some commonalities with addition but also introduces new complexes like borrowing. Extending our approach to handle such nuances would demonstrate its generalisability.
For multiplication, an interesting avenue is to first pre-train a large transformer model on addition using our verified modules as a starting component. Then, expand the model’s capacity and fine-tune on multiplication data. This could facilitate more rapid acquisition of the multiplication algorithm by providing a strong inductive bias grounded in robust addition skills. Specific experiments could then identify the emergence of new multiplication-specific subroutines and their integration with the addition circuits.
Furthermore, domains like symbolic AI, program synthesis, or general reasoning may benefit from embedding multiple specialized algorithmic components like our addition circuit within larger language models. Insights from our work could guide the controlled emergence and combination of diverse task-specific capabilities. Overall, while making models more interpretable is valuable, the ultimate aim is developing safer, more reliable, and more controllable AI systems. Our work highlight one path toward that goal through understanding of addition in neural network computations.
@@ 11 Reproducibility Statement
To facilitate the reproduction of our empirical results on understanding and interpreting addition in one-layer transformers, and further studying the properties of more complext transformers on more complex tasks that would build on a single layer, we release all our code and resources used in this work. Furthermore, we offer explicit constructions of one layer transformers used in this paper.
@@ Appendix A Appendix - Number of Attention Heads
The model can be successfully trained with 2, 3 or 4 attention heads (Refer Figs. 101112). However as described in Fig. 11, the 3 attention heads case is easier to interpret than the 2 or 4 heads cases.
Figure 10: For 5 digit addition, using 2 attention heads works, but the model attends to multiple token pairs in a single head, suggesting that multiple tasks are being packed into a single head, which makes it harder to interpret.
@@ Appendix B Appendix - Training Loss by Number of Digits
The model can be successfully trained with 5, 10, 15, etc digits (Refer Figs. 131415). For a given loss threshold, 15 digit addition takes longer to train than 10 digit addition, which takes longer to train than 5 digit addition.
Figure 13: The per-digit loss curves over 1500 training epochs for 5 digit addition.
Figure 14: The per-digit loss curves over 1500 training epochs for 10 digit addition.
Figure 15: The per-digit loss curves over 1500 training epochs for 15 digit addition.
@@ Appendix C Appendix - Attention Patterns by Number of Digits
The model can be successfully trained with 5, 10, 15, etc digits (Refer Figs. 161718). The attention patterns for these cases show similarities that aid interpretation.
Figure 16: For 5 digit addition, and 3 attention heads, the attention pattern has a strong double-staircase shape. Each step of the staircase is 3 blocks wide, showing the attention heads are attending to different input tokens in each row.
Figure 17: For 10 digit addition, and 3 attention heads, the attention pattern has a strong double-staircase shape. Each step of the staircase is 3 blocks wide, duplicating the Fig. 16 pattern.
Figure 18: For 15 digit addition, and 3 attention heads, the attention pattern has a strong double-staircase shape. Each step of the staircase is 3 blocks wide, duplicating the Fig. 16 pattern.
@@ Appendix D Appendix - Model Configuration
A Colab notebook was used for experimentation:
@@
  • The key parameters (which can all be altered) are:
@@
  • During a training run the model processes about 1.5 million training datums. For the 5 digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So, the training data is much less than 1% of the possible questions.
  • US9 cascades (e.g. 44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare. To speed up training, the data generator was enhanced to increase the frequency of these cases in the training data.
@@ Appendix E Appendix - Solving a jigsaw using many people
Understanding transformer algorithm implementation requires new ways of thinking.
Here is a useful analogy for changing how we approach a problem to more closely resemble how a Transformer would; a jigsaw puzzle. A single person could solve it using a combination of meta knowledge of the problem (placing edge pieces first), categorisation of resources (putting like-coloured pieces into piles), and an understanding of the expected outcome (looking at the picture on the box).
@@ Appendix F Appendix - Use Sum 9 Training Loss Graph
Fig. 19 shows a training loss curve for just BA and UC1 tasks. The model has high loss on Use Sum 9 cascades shown as high variability in the loss graph (See Fig. 20).
Figure 19: Training loss for digit 3 showing that, in Phase 2, the refining of Use Carry 1 (red line) lags Base Add (blue line). This supports the claim that Base Add and Use Carry 1 are learnt and refined separately and perform different calculation tasks.
Figure 20: Training loss for digit 3 for the Base Add, Use Carry 1 and Use Sum 9 tasks, showing the model struggles to reduce loss on Use Sum 9 (green line) compared to Base Add and Use Carry 1 .
@@ Appendix G Appendix - Model Algorithm as Pseudocode
The Algorithm 1 table below reproduces the model’s addition algorithm summarised in Fig. 9 as pseudocode. This code calculates the largest digit first, handling simple (but not cascading) UseSum9 cases, and returns the answer.
@@
3:
4:   for all $i=0,\dots,n\ digits-1$ do $\triangleright$ Loop over each digit
5:      $\text{pos}\leftarrow n\ digits-i-1$triangleright Current position from the right
@@
11:         if $q1_{\text{prev prev pos}}+q2_{\text{prev prev pos}}\geq 10$ then
@@
18:         if $q1_{\text{prev pos}}+q2_{\text{prev pos}}\geq 10$ then
@@
25:         if $q1_{\text{prev pos}}+q2_{\text{prev pos}}==9$ then
@@
40:      $\text{digitanswer}\leftarrow\textsc{MODULUS}(\text{digitanswer},10)$triangleright Correct the current digit if it’s $\geq 10$
41:      $\text{answer}\leftarrow\text{digitanswer}+\text{answer}\times 10$triangleright Concatenate the current digit to the answer
42:   end for
43:
44:   return answer $\triangleright$ Return the final calculated answer
45:end function
@@ Appendix H Appendix - The Loss function and loss measures
The final training loss varies with the number of digits in the question as shown in Tab. 1. The final training loss for each digit in the question varies as shown in Tab. 2.
@@
Table 2: For 5-digit addition, final training loss per digit. The more digits, the longer the model needs to train to a given level of loss.
We categorize each training question by which calculations its need to get the correct answer:
  • BA : BaseAdd questions only need “base add” calculations.
  • MC1 : MakeCarry1 questions need both “use carry 1” and “base add” calculations.
  • US9 : UseSum9 questions need “use sum 9”, “use carry 1” and “base add” calculations.
We measured the loss per question category as shown in Tab. 3. The frequency of these question types differ significantly ( BA = 61%, MC1 =33%, US9 =6% ) in the enriched training data so the final training loss values, at this level of detail, are not very informative.
@@
After training, we used the model to give answers to questions. To understand whether the calculations at position n are important, we look at the impact on loss of ablating all attention heads at that token (using the TransformerLens framework, zero ablation of the blocks.0.hook_resid_post data set, the above loss function, and a “cut off” loss threshold of 0.08). Tab. 4 shows sample results. The results shows that all important calculations are completed at just 6 of the 18 (question and answer) token positions.
@@
Table 4: For 5 digit addition, for the 18 tokens / calculation rows, how ablating each token/row impacts the calculation loss.
For deeper investigation, we created 100+ hand-curated test questions. They cover all question types (BA, MC1 and MS9) and all the answer digits (A5 .. A0) that these question types can occur in. These test cases were not used in model training. Example test questions include:
@@
 11 Nyyyyy: 47, NyNyyy: 5, yNyyyy: 7, yyNyyy: 7, yyyNyy: 1, …
12 yNyyyy: 97, yNNyyy: 7, NNyyyy: 3, Nyyyyy: 2, yyyNyy: 1, …
13 yyNyyy: 85, NyNyyy: 3, yNNyyy: 2, Nyyyyy: 3, yNyyyy: 5, …
14 yyyNyy: 72, yyNNyy: 4, NyyNyy: 3, yNyNyy: 3, Nyyyyy: 2, …
15 yyyyNy: 74, yyNyNy: 3, Nyyyyy: 2, NyyyNy: 3, yNyyNy: 3, …
16 yyyyyN: 82, yyNyyN: 2, NyyyyN: 4, yNyyyN: 4, yyNyyy: 9, …
Table 5: For 5 digit addition, ablating all the attention heads in each row (aka step) impacts answer correctness in a regular pattern. Here, an ‘N’ means that answer digit was incorrect in the predicted answer. In each step, the failure count for the top grouping is > 6 times the next most common grouping - evidence that each step calculates one answer digit. From row 11, each row mainly impacts one answer digit - in the same order as the model predicts answer digits.
@@ 11 Nyyyyy: 47, NyNyyy: 5, yNyyyy: 7, yyNyyy: 7, yyyNyy: 1, …
Appendix I Appendix - Algorithm Re-use
We explored whether the algorithm is re-used in other models by training a separate 1-layer 5-digit addition model with a different seed, optimiser. It also required the model to predict “+" as the first answer token as shown in Fig. 3.
Figure 21: We refer to individual tokens in a 5-digit addition question as D4, .. D0, and D’4, .., D’0 and the answer tokens as A6, .., A0. Note that A6 is always the “+" token.
Figs. 24,  25 and  23 show aspects of the behavior of this model. Fig. 26 shows the nodes in the model where the node purpose can be identified by automated search using behavior filtering and ablation intervention testing.
Figure 22: This shows the token positions where questions fail when we ablate each node in the 5-digit 1-layer 3-head addition model. The model only uses nodes in token positions P12 to P17 (answer digits A6 to A1).
Figure 23: This map shows the percentage of enriched questions that fail when we ablate each node in the 5-digit 1-layer 3-head addition model. The model only uses nodes in token positions P12 to P7 (answer digits A6 to A1). Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not used by the model.
Figure 24: This map shows the input tokens each attention head attends to at each token position in the new model. The new model only uses attention heads in token positions P12 to P17 (answer tokens A6 to A1). Each head attends to a pair of input tokens. Each column attends to three different input token positions. These behaviors are the same in the new and previous models.
Figure 25: This map shows the answer digit(s) A0 .. A5 impacted when we ablate each attention head and MLP layer in the new model. This behavior is the same in the new and previous models. Note that A6 is the constant “+" answer token.
Figure 26: This map shows the calculation tasks BA, MC1 and US9 located in this model by automated search using behavior filtering and ablation intervention testing. For nodes containing “??" the automated search did not identify the purpose of the node.